State whether the following statements are true or false. Justify yourself with the help of
a short proof or a counter example(1)Equation x^2 ( y − px) = yp^2 is reducible to clairaut’s form
1
Expert's answer
2020-03-23T15:14:58-0400

This statement is false. The Clairaut equation is "y=xp+f(p)" where "p=y'".

"x^2(y-px)=yp^2" is reducible to form "y=xp+\frac{y}{x^2}f(p)"